# Bias Behind the Wheel: Fairness Analysis of Autonomous Driving Systems

### v3

@@ 5.2.1. Implication for researchers
Our empirical results offer valuable research opportunities for researchers. 1) Fairness improvement of autonomous driving systems via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific post-processing image editing techniques to adjust contrast and brightness of captured input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Real-time adaptive fairness improvement of autonomous driving systems. Adapting software systems to increasingly dynamic environments poses significant challenges in software development, an area that has long been a important task in SE research (Qureshi and Perini, 2009). Our empirical findings contribute to this ongoing task by introducing a new challenge: designing real-time adaptive fairness improvement methods for autonomous driving systems. Specifically, our research reveals that autonomous driving systems exhibit greater biases under specific driving scenarios influenced by changes in brightness and contrast. Given the common occurrence of such changes in real-world driving conditions, there is an urgent need to develop specific methods to ensure the consistent fairness of autonomous driving systems in response to these environmental dynamics. 3) Fairness improvement of autonomous driving systems via multi-objective optimization. Given our finding that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. 4) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 5) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage fellow researchers to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
@@ 5.3. Threats to Validity
Selection of pedestrian detectors. Our study is based on eight pedestrian detectors, which may lead to possible selection bias in this study. To mitigate this threat, we select representative pedestrian detectors based on two considerations. On one hand, the selected detectors are widely-studied in pedestrian detection and autonomous driving literature (Cao et al., 2020; Zhang et al., 2016, 2017; Yu et al., 2020). On the other hand, the selected models cover the two typical types of pedestrian detection methods (i.e., one-stage detection and two-stage detection), ensuring a comprehensive representation of the techniques used in the field.
Selection of datasets. The choice of datasets presents a potential threat to validity. To address this concern, we use four real-world datasets extensively explored in autonomous driving research, including a total of 8,311 images. We augment the datasets with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. The extensive scale of data ensures the reliability of our results.
@@ 6. Conclusion
This paper presents the first comprehensive study on fairness of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors regarding gender, age, and skin tone. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. Additionally, pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Regarding skin tone, we observe balanced detection performance for light-skin and dark-skin individuals overall. As part of our contribution, we publicly release large-scale real-world pedestrian detection datasets with gender, age, and skin tone labels. These datasets aim to facilitate future fairness research in autonomous driving. The insights gained in this study can pave the way for more fair and unbiased autonomous driving systems in the future.

### v4

@@ 5.3.1. Implication for researchers
Our empirical results offer valuable research opportunities for researchers. 1) Fairness improvement of autonomous driving systems via image editing. Our findings reveal that brightness and contrast of images can impact the fairness of pedestrian detection. Thus, to improve fairness, a practical solution is to design specific post-processing image editing techniques to adjust contrast and brightness of captured input images, such as increasing brightness and contrast levels to dynamically counterbalance existing bias towards children and female pedestrians in low-brightness and low-contrast conditions. Moreover, previous discussions in Section 5.1 show that these solutions also have the potential to improve detection performance. 2) Real-time adaptive fairness improvement of autonomous driving systems. Adapting software systems to increasingly dynamic environments poses significant challenges in software development, an area that has been an important task in SE research (Qureshi and Perini, 2009). Our empirical findings contribute to this ongoing task by introducing a new challenge: designing real-time adaptive fairness improvement methods for autonomous driving systems. Specifically, our research reveals that autonomous driving systems exhibit greater biases under specific driving scenarios influenced by changes in brightness and contrast. Given the common occurrence of such changes in real-world driving conditions, there is an urgent need to develop specific methods to ensure the consistent fairness of autonomous driving systems in response to these environmental dynamics. 3) Fairness improvement of autonomous driving systems via multi-objective optimization. Given our finding that the fairness-performance trade-off may not hold in the context of pedestrian detection, researchers have the opportunity to develop multi-objective optimization strategies for model training, enabling the concurrent optimization of fairness and detection performance. At the same time, based on previous findings regarding brightness and contrast, a possible approach could involve integrating fairness constraints during model training by incorporating contrast-aware or brightness-aware regularization techniques. The model could penalize overly confident predictions in low-contrast or low-brightness conditions, encouraging it to treat predictions more cautiously in these scenarios. Additionally, augmenting data by enhancing these attributes in images of underrepresented groups (e.g., children in low-brightness conditions) could help the model generalize better. Synthetic image generation that varies brightness and contrast during training may further expose the model to diverse brightness and contrast conditions, potentially reducing bias. 4) Addressing inherent limitations for detecting small objects of pedestrian detectors. For age bias, it is important to address the inherent limitations of pedestrian detectors, particularly their difficulty in accurately detecting small pedestrians. These implications underline the importance of factoring these elements into effective fairness improvement strategies. 5) Open datasets and annotations for fairness assessment. There is a shortage of available datasets annotated with a wide range of demographic labels for pedestrian detection. To address this constraint, we have undertaken the task of manually annotating four datasets with information related to gender, age, and skin tone, and publicly released these annotated datasets. It is worth noting that there are additional demographic attributes that also need to be considered in the context of fairness, such as disability status and religion. We encourage researchers and practitioners to join us in the effort to promote fairness in pedestrian detection by creating and sharing datasets that encompass various demographic attributes.
@@ 5.4. Threats to Validity
Selection of pedestrian detectors. Our study is based on eight pedestrian detectors, which may lead to possible bias in this study. To mitigate this threat, we select representative pedestrian detectors based on two considerations. On one hand, the selected detectors are widely-studied in pedestrian detection and autonomous driving literature (Cao et al., 2020; Zhang et al., 2016, 2017; Yu et al., 2020). On the other hand, the selected models cover the two typical types of pedestrian detection methods (i.e., one-stage detection and two-stage detection), ensuring a comprehensive representation of the techniques used in the field.
Selection of datasets. The choice of datasets presents a potential threat to validity. Current literature on autonomous driving sometimes relies on automated techniques to simulate adverse weather conditions such as fog, rain, and snow. However, these synthetic images may not accurately reflect real-life conditions (Zhang et al., 2022) (e.g., simply simulating fog by applying a set of filters over the original image, without considering the complex transformations that occur in actual road scenes under such conditions). Although real-world datasets for adverse weather like fog now exist (Bijelic et al., 2020), accurately annotating demographic attributes of pedestrians under foggy conditions remains challenging in the context of this study. These weather effects can obscure important features such as skin tone and blur physical outlines, making it difficult to discern attributes like gender and age. To address these concerns, we utilize four real-world datasets extensively explored in autonomous driving research, comprising a total of 8,311 images. Additionally, we have augmented these datasets with 16,070 gender labels, 20,115 age labels, and 3,513 skin tone labels. The extensive scale of data ensures the reliability of our results and mitigates the limitations associated with synthetic images.
@@
Selection of sensitive attributes. The selection of sensitive attributes may introduce potential limitations to our study. To mitigate this concern, we focus on attributes that are most identifiable in autonomous driving images and are recognized as the three most extensively studied sensitive attributes in the fairness literature: gender, age, and skin tone (Hort et al., 2024). While other sensitive attributes, such as disability status, precise age determinations for young and elderly individuals, or nationality, are relevant to fairness considerations, they present significant challenges for consistent identification from typical vehicle-mounted camera data due to ambiguous feature judgment. In contrast, gender, age (adult and child), and skin tone are more readily observable characteristics in such images, allowing for more reliable annotations.
@@ 6. Conclusion
This paper presents the first comprehensive study on fairness testing of eight state-of-the-art pedestrian detectors, using four widely-studied testing datasets. We investigate the fairness aspects of these detectors regarding gender, age, and skin tone. Furthermore, we conduct an in-depth analysis of fairness in various driving scenarios. Our findings reveal significant bias in the current pedestrian detectors, particularly towards children. Additionally, pedestrian detectors demonstrate significant gender biases during night time, potentially exacerbating the prevalent societal issue of female safety concerns during nighttime out. Regarding skin tone, we observe balanced detection performance for light-skin and dark-skin individuals overall. As part of our contribution, we publicly release large-scale real-world pedestrian detection datasets with gender, age, and skin tone labels. These datasets aim to facilitate the future fairness research in autonomous driving. The insights gained in this study can pave the way for more fair and unbiased autonomous driving systems in the future.
@@ 7. Data Availability
Acknowledgement
This work was supported by the National Natural Science Foundation of China under the grant number 62325201, as well as by the Center for Data Space Technology and Systems at Peking University. Xinyue Li, Ying Zhang, and Xuanzhe Liu are also affiliated with the Key Lab of High Confidence Software Technologies (Peking University), Ministry of Education, Beijing, China.